OpenStudy (anonymous):

How many days and hours are in 7500 mins?

OpenStudy (anonymous):

lol

OpenStudy (hoblos):

each 60min = 1hr each 24hr = 1 day 7500 min = 7500/60 = 125 hr 125/24hr = 5 days and 5hrs

OpenStudy (anonymous):

5 days and 5 hours